# Textual semantics and machine learning methods for data product pricing

### v1

@@ Abstract
Keywords: Data product price, Machine learning, Textual representation, Feature importance analysis
1 Introduction
In the era of digital economy, data referred to as the “new oil” has become an invaluable resource (Pei,, 2020; Zhu et al.,, 2024). Governments, enterprises, and individuals worldwide are increasingly leveraging data-driven approaches to unlock substantial economic value. For example, the U.S. healthcare sector could generate more than USD 300 billion in value annually by fully harnessing big data.111Big data: The next frontier for innovation, competition, and productivity, McKinsey Global Institute, https://www.mckinsey.com/capabilities/tech-and-ai/our-insights/big-data-the-next-frontier-for-innovation In the retail sector, the adoption of dynamic pricing algorithms based on consumer behavioral data has been shown to increase revenue by 11% to 19% (Fisher and Raman,, 2018). Similarly, the UK’s HM Revenue & Customs (HMRC) recovered GBP 1.4 billion in unpaid taxes in a single year through its Connect system, which integrates 28 data sources to support anomaly detection and risk modeling (Maciejewski,, 2017). The rapid growth in data-driven demand has fueled the rise of Data Marketplaces, which is an online platforms that bring together data providers, consumers, and brokers to facilitate the exchange of data products (Zhang et al.,, 2023). According to a report by Grand View Research (2025), the global DMs market was valued at approximately USD 1.49 billion in 2024 and is predicted to reach USD 5.73 billion by 2030.222Data Marketplace Platform Market (2025-2030), Grand View Research, https://www.grandviewresearch.com/industry-analysis/data-marketplace-market-report
@@ 1 Introduction
Although existing studies have made initial investigation in applying machine learning methods to data product pricing, most current work still relies primarily on numerical features (e.g., dataset size, update frequency, category) to characterize data products. Few studies extracts value-relevant information from textual semantics. In practice, on data marketplaces such as AWS Data Exchange and Datarade, data products are typically presented through natural language descriptions, including data product titles, summaries, detailed descriptions, and application scenarios. The semantic content embedded in this text not only reflects the product’s thematic and domain attributes but also conveys implicit signals about its market positioning and perceived value (Gan et al.,, 2025). Although Zhu et al., (2024) made an early attempt to incorporate textual features into data product pricing, their approach relied solely on a single mBERT-based textual representation. It did not consider other widely used textual representations such as Bag-of-Words (BOW), TF-IDF, or topic-based models. Moreover, data product pricing commonly involves two forms, i.e., precise (continuous) pricing and tiered (categorical) pricing, yet their study did not investigate how different textual representation methods may influence these two types of pricing tasks.
@@
To investigate how different textual feature representation methods influence various types of data product pricing tasks, we conduct a systematic empirical analysis. First, we extract textual information from data products and represent it using five prevalent textual feature engineering methods: Bag-of-Words, TF-IDF, Word2Vec, LDA, and BERTopic. These methods capture textual semantics from multiple perspectives and provide a comprehensive quantification of data product descriptions. Second, we integrate these textual features with structured attributes such as numerical and categorical variables, and construct pricing models using a diverse set of machine learning algorithms, including linear regression, neural networks, decision trees, support vector regression, random forests, and XGBoost. In the empirical analysis, we design two prediction tasks, a regression task for modeling the continuous prices of data products and a classification task that categorizes prices into discrete tiers, to thoroughly compare the performance of different textual representation methods across distinct prediction scenarios. Finally, we conduct feature-importance analysis using mRMR for feature selection and SHAP for model interpretability. These analyses identify the key features of data product pricing and further validate the critical role of textual semantic features in improving pricing prediction performance.
@@ 2.2 Machine learning approaches to data pricing
To capture more complex, non-linear relationships in data pricing, many studies have adopted powerful ensemble learning models known for their high predictive accuracy. For instance, Zhao, (2024) uses an XGBoost model for prediction, pairing it with SHAP (SHapley Additive exPlanations) to maintain high accuracy while ensuring model interpretability. For pricing high-dimensional and non-linear data, such as in healthcare, Shang et al., (2025) leverage the capabilities of LightGBM. A distinct trend in this area involves creating novel ensemble strategies by combining multiple heterogeneous models. Lv, (2025) demonstrates that a stacking ensemble, which combines three base models and a fusion model, outperforms any individual model. Hao et al., (2023) are among the first to apply a Rank-Prune-Average strategy, which averages predictions from the best-performing models in a heterogeneous set. Building on this, Hao et al., (2024) develop a more sophisticated fusion method that uses Random Forest (RF) for feature selection, K-means for clustering, and a ”gravity-based” weighting strategy to combine fifteen different models.
@@ 3 Methodology
This section mainly consists of four components: feature representation, feature selection models, data price prediction models, and performance evaluation metrics.
@@ 3.1.1 Text feature representation
where is the feature mapping of document , the BOW vector corresponding to document , is the frequency of the vocabulary term, is the multiset of tokens obtained after tokenizing document , is an indicator function, which equals 1 if the token is equal to the vocabulary word , and 0 otherwise.
@@ 3.1.2 Numerical and categorical feature representation
The provider name is represented as a binary variable (0-1): if the provider is a publicly listed company, the value is set to 1; otherwise, it is 0.
@@
Refund policy describes the refund terms, categorized into five levels. For this feature, the classification was performed using DeepSeek, based on the textual descriptions of refund policies, with the detailed prompt instructions provided in the Appendix.
@@
For data product industry, we employ DeepSeek (model version: DeepSeek-V3.2-Exp) to parse the product information and classify the products into 12 industries. The industry taxonomy is based on Amazon’s organizational structure of data products, while making several domain-specific adjustments: for instance, agriculture and energy are separated as independent resource-based categories, and marketing data is further divided into e-commerce and offline retail scenarios. The specific industry classification scheme and corresponding prompt design are presented in the Appendix.
@@ 3.4 Performance evaluation
The dataset is divided using 5-fold cross-validation. For the regression task, the evaluation metrics included Mean Squared Error (MSE), Root Mean Squared Error (RMSE), and Mean Absolute Percentage Error (MAPE). The calculation formulas are as follows:
where denotes the predicted value and represents the actual value.
@@ 4.1 Data description
In this study, we collect data from Amazon, excluding data products with free prices, and primarily focuses on subscription-based data products. AWS Data Exchange is a comprehensive marketplace for third-party datasets. It provides a secure and compliant environment, seamlessly integrates with AWS and third-party tools and services, and supports unified billing and subscription management. The features of the dataset include product name, product detail, product description, provider name, volume, history, data sample, support email, support URL, and refund policy, as summarized in Table 1.
@@
We can see from Table 2 that in term of the feature (Data provider is listed), the mean value is only 0.08, indicating that the majority of data providers are non-listed enterprises or institutions, with only a small proportion being publicly listed companies. The Data volume number has a mean of 1.74, a standard deviation of 4.95, and a maximum value of 89, suggesting substantial variation in dataset size across different data products. For Sensitive level, the mean value is 0.25, implying that most data products do not contain sensitive information. Regarding version-related variables, the mean value of Future version is 0.94, indicating that most products support future updates, while the mean of Historical version is 1.68, showing that a considerable number of products provide access to historical data.
@@
The distribution of different industries is presented in Table 3.
@@
Overall, data products are mainly concentrated in three primary categories: E-commerce and Business Data, Financial Services, and Retail and Location Data, with counts of 548, 469, and 459 respectively, representing the largest proportions in the dataset. Following these are Public Sector Data (300) and Healthcare and Life Sciences Data (242), indicating notable representation in government and medical domains. Media and Entertainment Data (234) and Cars and Automotive Data (225) fall into the medium range, reflecting moderate activity in these sectors. In contrast, Environmental Data (208), Gaming Data (201), Telecommunications Data (183), Resources Data (160), and Manufacturing Data (147) account for smaller shares.
@@ 4.2 Regression results comparison based on different feature representation methods
To compare the predictive performance of different text feature representation methods (BOW, TF-IDF, Word2Vec, LDA, and BERTopic) on data-product pricing, multiple prediction models are employed, including Linear Regression (LR), Artificial Neural Network (ANN), Decision Tree Regression (DT), Support Vector Regression (SVR), Random Forest (RF), and XGBoost. The prediction results are shown in Table 4 where ME represents the mean error across the six predictive models, Rank indicates the ranking results based on the ME values, and the values in bold font indicate the best results among different text representation methods under the same prediction model.
| Method | LR | ANN | DT | SVR | RF | XGBoost | ME | Rank |
@@
| RMSE | ||||||||
| BOW | 2.4938 | 1.1564 | 1.2826 | 1.2243 | 0.9359 | 1.0950 | 1.3647 | 4 |
| TFIDF | 1.3434 | 1.1551 | 1.4098 | 1.2161 | 1.0174 | 1.1202 | 1.2103 | 2 |
| Word2vec | 1.2338 | 1.1713 | 1.4552 | 1.2205 | 1.0177 | 1.0765 | 1.1958 | 1 |
| LDA | 1.4958 | 1.3311 | 1.5059 | 1.4269 | 1.1226 | 1.2408 | 1.3539 | 3 |
| BERTopic | 1.6474 | 1.5132 | 1.7090 | 1.5533 | 1.2731 | 1.3913 | 1.5146 | 5 |
@@
Based on the results in Table 4, we first find that different text feature representation methods exhibit varying predictive performances. Word2Vec achieves the best overall results, ranking first in both MSE and RMSE, indicating that its semantic representation provides the highest predictive capability for data-product pricing. TF-IDF follows closely, ranking first in MAPE, which suggests that traditional statistical methods remain robust when dealing with high-dimensional and sparse text data. In contrast, BERTopic performs the worst, likely due to insufficient data volume or overly coarse clustering granularity, leading to poor generalization of its topic representations. Second, regarding the results of different predictive models: LR performs relatively poorly across all feature representation methods, while ensemble models such as RF and XGBoost achieve better results. This indicates that the relationship between data-product prices and their textual features is inherently nonlinear.

### v2

@@ Abstract
[…]
Data product price, Machine learning, Textual representation, Feature importance analysis1 Introduction
In the era of digital economy, data referred to as the “new oil” has become an invaluable resource (Pei,, 2020; Zhu et al.,, 2024). Governments, enterprises, and individuals worldwide are increasingly leveraging data-driven approaches to unlock substantial economic value. For example, the U.S. healthcare sector could generate more than USD 300 billion in value annually by fully harnessing big data.111Big data: The next frontier for innovation, competition, and productivity, McKinsey Global Institute, https://www.mckinsey.com/capabilities/tech-and-ai/our-insights/big-data-the-next-frontier-for-innovation. In the retail sector, the adoption of dynamic pricing algorithms based on consumer behavioral data has been shown to increase revenue by 11% to 19% (Fisher and Raman,, 2018). Similarly, the UK’s HM Revenue & Customs (HMRC) recovered GBP 1.4 billion in unpaid taxes in a single year through its Connect system, which integrates 28 data sources to support anomaly detection and risk modeling (Maciejewski,, 2017). The rapid growth in data-driven demand has fueled the rise of Data Marketplaces, which is an online platforms that bring together data providers, consumers, and brokers to facilitate the exchange of data products (Zhang et al.,, 2023). According to a report by Grand View Research (2025), the global DMs market was valued at approximately USD 1.49 billion in 2024 and is predicted to reach USD 5.73 billion by 2030.222Data Marketplace Platform Market (2025-2030), Grand View Research, https://www.grandviewresearch.com/industry-analysis/data-marketplace-market-report.
@@
Although existing studies have made initial investigation in applying machine learning methods to data product pricing, most current work still relies primarily on numerical features (e.g., dataset size, update frequency, category) to characterize data products. Few studies extracts value-relevant information from textual semantics. In practice, on data marketplaces such as Amazon Data Exchange and Datarade, data products are typically presented through natural language descriptions, including data product titles, summaries, detailed descriptions, and application scenarios. The semantic content embedded in this text not only reflects the product’s thematic and domain attributes but also conveys implicit signals about its market positioning and perceived value (Gan et al.,, 2025). Although Zhu et al., (2024) made an early attempt to incorporate textual features into data product pricing, their approach relied solely on a single mBERT-based textual representation. It did not consider other widely used textual representations such as Bag-of-Words (BOW), TF-IDF, or topic-based models.
Current challenges in using textual data for data product pricing include the following aspects. First, there exist multiple approaches to text representation, ranging from context-independent methods such as BOW and TF-IDF, to context-aware methods such as Word2Vec, as well as topic-model-based representations. However, it remains unclear which types of textual representations are most appropriate for characterizing data products and for supporting pricing decisions. Second, data product pricing typically involves two main paradigms: tiered pricing and precise pricing. Tiered pricing enables rapid classification of data products into different price levels, thereby providing an approximate price range, while precise pricing focuses on determining an explicit price to serve as a basis for negotiation between data providers and data consumers. The respective roles and effectiveness of different text representation methods in these distinct pricing tasks have not yet been systematically investigated.
@@
Therefore, to investigate how different textual feature representation methods influence various types of data product pricing tasks, we conduct a systematic empirical analysis. First, we extract textual information from data products and represent it using five prevalent textual feature engineering methods: Bag-of-Words, TF-IDF, Word2Vec, LDA, and BERTopic. These methods capture textual semantics from multiple perspectives and provide a comprehensive quantification of data product descriptions. Second, we integrate these textual features with structured attributes such as numerical and categorical variables, and construct pricing models using a diverse set of machine learning algorithms, including linear regression, neural networks, decision trees, support vector regression, random forests, and XGBoost. In the empirical analysis, we design two prediction tasks, a regression task for modeling the continuous prices of data products and a classification task that categorizes prices into discrete tiers, to thoroughly compare the performance of different textual representation methods across distinct prediction scenarios. Finally, we conduct feature-importance analysis using mRMR for feature selection and SHAP for model interpretability. These analyses identify the key features of data product pricing and further validate the critical role of textual semantic features in improving pricing prediction performance.
@@ 2.2 Machine learning approaches to data pricing
To capture more complex, non-linear relationships in data pricing, many studies have adopted powerful ensemble learning models known for their high predictive accuracy. For instance, Zhao, (2024) uses an XGBoost model for prediction, pairing it with SHAP (SHapley Additive exPlanations) to maintain high accuracy while ensuring model interpretability. For pricing high-dimensional and non-linear data, such as in healthcare, Shang et al., (2025) leverage the capabilities of LightGBM. A distinct trend in this area involves creating novel ensemble strategies by combining multiple heterogeneous models. Lv, (2025) demonstrates that a stacking ensemble, which combines three base models and a fusion model, outperforms any individual model. Hao et al., (2023) are among the first to apply a Rank-Prune-Average strategy, which averages predictions from the best-performing models in a heterogeneous set. Building on this, Hao et al., (2024) develop a more sophisticated fusion method that uses Random Forest (RF) for feature selection, K-means for clustering, and a “gravity-based” weighting strategy to combine fifteen different models.
@@ 3 Methodology
This section mainly consists of problem statement, feature representation, feature selection models, data price prediction models, and performance evaluation metrics.
3.1 Problem statement
In this study, the model inputs primarily consist of three types of features derived from data products: textual descriptions, numerical attributes, and categorical attributes. For textual data, we employ multiple representation methods, including Bag-of-Words (BOW), Term Frequency–Inverse Document Frequency (TF-IDF), Word2Vec, Latent Dirichlet Allocation (LDA), and BERTopic. For categorical data, we adopt one-hot encoding to transform categorical attributes into numerical feature vectors. The modeling framework primarily relies on a range of machine learning methods, including Logistic Regression (LR), Artificial Neural Networks (ANN), Decision Trees (DT), Support Vector Machines (SVM), Random Forests (RF), and XGBoost.
The model outputs involve two tasks: a price-interval classification task and a point-price regression task. We include price-interval prediction as a complementary task to point-price regression for both economic and practical reasons. First, data are information goods whose realized value is buyer-specific: different buyers combine the same dataset with different existing data assets, models, and use cases, leading to substantial dispersion in willingness to pay (Farboodi et al.,, 2025). In such settings, a predicted range can be more actionable for screening and negotiation than a single point estimate. Second, data marketplaces commonly implement “versioning” or tiered pricing and may adjust prices dynamically as more information about demand, usage, and product quality becomes available (Pei,, 2020). Our interval prediction therefore aligns with a realistic workflow in which products are first positioned into coarse price tiers (e.g., for rapid listing or benchmarking) and then refined into an exact quote when needed. Finally, given the highly right-skewed, long-tailed distribution of data-product prices, tier prediction provides a more robust and interpretable reference that is less sensitive to extreme values, while regression results offer finer granularity for applications requiring precise pricing.
@@ 3.2.1 Text feature representation
where denotes the feature mapping function, is the corresponding the BOW vector corresponding to document . The element represents the frequency of the vocabulary term in document . is the multiset of tokens obtained after tokenizing document , is an indicator function, which equals 1 if the token is equal to the vocabulary word , and 0 otherwise.
@@ 3.2.2 Numerical and categorical feature representation
The provider name is represented as a binary variable (0–1): if the provider is a publicly listed company, the value is set to 1; otherwise, it is 0.
@@
Refund policy describes the refund terms, categorized into five levels. For this feature, the classification was performed using DeepSeek, based on the textual descriptions of refund policies, with the detailed prompt instructions provided in the Appen-dix.
@@
For data product industry, we employ DeepSeek (model version: DeepSeek-V3.2-Exp) to parse the product information and classify the products into 12 industries.333The industry classification of data products was conducted using the DeepSeek accessed via direct API calls. This choice was motivated by the reason that the classification task relies primarily on semantic understanding of textual descriptions rather than domain-specific labeled training data, which makes general-purpose large language model appropriate. The industry taxonomy is based on Amazon’s organizational structure of data products, while making several domain-specific adjustments: for instance, agriculture and energy are separated as independent resource-based categories, and marketing data is further divided into e-commerce and offline retail scenarios. The specific industry classification scheme and corresponding prompt design are presented in the Appendix.
@@ 3.5 Performance evaluation
The dataset is divided using 5-fold cross-validation. For the regression task, the evaluation metrics included Mean Squared Error (MSE), the coefficient of determination () and Mean Absolute Percentage Error (MAPE). The calculation formulas are as follows:
where denotes the predicted value, represents the actual value and .
@@ 4.1 Data description
4.1.1 Sample
In this study, we collect data from Amazon Data Exchange, which is a comprehensive marketplace for third-party datasets. It provides a secure and compliant environment, seamlessly integrates with Amazon Web Services and third-party tools and services, and supports unified billing and subscription management. We exclude data products with free prices, and primarily focuses on subscription-based data products. As a result, the sample consists of 2,073 data products.
4.1.2 Feature
The features of the dataset include product name, product detail, product description, provider name, volume, history, data sample, support email, support URL, and refund policy, as summarized in Table 1. The descriptive statistics of the features are presented in Table 2. We can see from Table 2 that in term of the feature (Data provider is listed), the mean value is only 0.08, indicating that the majority of data providers are non-listed enterprises or institutions, with only a small proportion being publicly listed companies. The Data volume number has a mean of 1.74, a standard deviation of 4.95, and a maximum value of 89, suggesting substantial variation in dataset size across different data products. For Sensitive level, the mean value is 0.25, implying that most data products do not contain sensitive information. Regarding version-related variables, the mean value of Future version is 0.94, indicating that most products support future updates, while the mean of Historical version is 1.68, showing that a considerable number of products provide access to historical data.
@@ 4.1.2 Feature
The distribution of different industries is presented in Table 3. To evaluate the reliability of the classification, we conducted a manual validation exercise. Specifically, we drawed a random sample of 120 data products from the full dataset, covering all 12 industry categories, with 10 observations per category. Then we reviewed and assigned each sample to an industry label based on product’s textual title and description. The manually assigned labels were then compared with the generated labels to assess accuracy. Validation results indicate a high level of classification accuracy of 0.94 (113 correct classifications out of 120). These results suggest that misclassifications are limited and do not materially affect the overall industry distribution patterns. Manual validation results of industry classification accuracy are reported in Table 4.
@@
| Industry | Correct | Total | Accuracy |
| E-commerce and Business Data | 10 | 10 | 1.00 |
| Financial Services | 10 | 10 | 1.00 |
| Retail and Location Data | 10 | 10 | 1.00 |
| Public Sector Data | 9 | 10 | 1.00 |
| Healthcare and Life Sciences Data | 10 | 10 | 1.00 |
| Media and Entertainment Data | 10 | 10 | 1.00 |
| Cars and Automotive Data | 9 | 10 | 0.90 |
| Environmental Data | 9 | 10 | 0.90 |
| Gaming Data | 8 | 10 | 0.80 |
| Telecommunications Data | 10 | 10 | 1.00 |
| Resources Data | 9 | 10 | 0.90 |
| Manufacturing Data | 9 | 10 | 0.90 |
| Total | 113 | 120 | 0.94 |
Overall, data products are mainly concentrated in three primary categories: E-comme-rce and Business Data, Financial Services, and Retail and Location Data, with counts of 548, 469, and 459 respectively, representing the largest proportions in the dataset. Following these are Public Sector Data (300) and Healthcare and Life Sciences Data (242), indicating notable representation in government and medical domains. Media and Entertainment Data (234) and Cars and Automotive Data (225) fall into the medium range, reflecting moderate activity in these sectors. In contrast, Environmental Data (208), Gaming Data (201), Telecommunications Data (183), Resources Data (160), and Manufacturing Data (147) account for smaller shares. It is worth noting that a single data product may be classified into multiple industries. Therefore, the total number reported in Table 3 exceeds the number of data products in the sample.
4.1.3 Data product price
@@
In addition to predicting the exact prices of data products, we also conduct price interval prediction. Specifically, the prices are divided into four quartiles, resulting in five distinct ranges from high to low. After categorizing the prices, the corresponding price ranges are defined as follows: Class 0: 208.33; Class 1: 212.5-416.67; Class 2: 418.5-1,250.0; Class 3: 1,258.33-3,175.0; Class 4: 3,200.0. Such interval-based prediction can provide data providers and data consumers with an approximate price range for reference.
@@ 4.2 Regression results comparison based on different feature representation methods
To compare the predictive performance of different text feature representation methods (BOW, TF-IDF, Word2Vec, LDA, and BERTopic) on data-product pricing, multiple prediction models are employed, including Linear Regression (LR), Artificial Neural Network (ANN), Decision Tree Regression (DT), Support Vector Regression (SVR), Random Forest (RF), and XGBoost. The prediction results are shown in Table 5 where MV represents the mean values across the six predictive models, Rank indicates the ranking results based on the MV value, and the values in bold font indicate the best results among different text representation methods under the same prediction model.
| Method | LR | ANN | DT | SVR | RF | XGBoost | MV | Rank |
@@
| (0.0871) | (0.2036) | (0.2367) | (0.0777) | (0.1539) | (0.1196) | (0.4820) | ||
| BOW | -1.1489 | 0.5703 | 0.4718 | 0.5204 | 0.7182 | 0.6157 | 0.2913 | 4 |
| (1.3601) | (0.0568) | (0.0581) | (0.0314) | (0.044) | (0.0371) | (0.7106) | ||
| TFIDF | 0.4221 | 0.5725 | 0.3614 | 0.5270 | 0.6684 | 0.5977 | 0.5249 | 2 |
| (0.061) | (0.0424) | (0.0887) | (0.0331) | (0.0341) | (0.0446) | (0.1144) | ||
| Word2Vec | 0.5106 | 0.5569 | 0.3202 | 0.5231 | 0.6679 | 0.6286 | 0.5346 | 1 |
| (0.0682) | (0.0828) | (0.0806) | (0.0369) | (0.0368) | (0.0386) | (0.1215) | ||
| LDA | 0.2845 | 0.4321 | 0.2704 | 0.3487 | 0.5962 | 0.5073 | 0.4065 | 3 |
| (0.0212) | (0.0522) | (0.1044) | (0.033) | (0.0373) | (0.0316) | (0.1293) | ||
| BERTopic | 0.1323 | 0.2673 | 0.0654 | 0.2284 | 0.4813 | 0.3811 | 0.2593 | 5 |
| (0.0102) | (0.0519) | (0.0629) | (0.0175) | (0.0391) | (0.0231) | (0.1542) | ||
@@
| (0.1096) | (0.0528) | (0.1112) | (0.1088) | (0.0481) | (0.0763) | (0.1519) | ||
Based on the results in Table 5, we first find that different text feature representation methods exhibit varying predictive performances. Word2Vec achieves the best overall results, ranking first in both MSE and , indicating that its semantic representation provides the highest predictive capability for data-product pricing. TF-IDF follows closely, ranking first in MAPE, which suggests that traditional statistical methods remain robust when dealing with high-dimensional and sparse text data. In contrast, BERTopic performs the worst, likely due to insufficient data volume or overly coarse clustering granularity, leading to poor generalization of its topic representations. Second, regarding the results of different predictive models: LR performs relatively poorly across all feature representation methods, while ensemble models such as RF and XGBoost achieve better results. This indicates that the relationship between data-product prices and their textual features is inherently nonlinear.